OpenStudy (anonymous):

The cost of 4 scarfs and 6 hats is 52.00.the cost of 2 hats is one dollar more than the cost of one scarf. what is the cost of one scarf? A.$4 B.$5 C.$6 D.$7

OpenStudy (jdoe0001):

\(\bf hats=h\qquad scarves=s\\ \quad \\ 4s+6h=52\qquad then\qquad h=s+1\qquad so\\ \quad \\ 4s+6(s+1) = 52\) solve for "s"

OpenStudy (jdoe0001):

hmmm one sec.... cost of 2 hats... ok... lemme rewrite that

OpenStudy (jdoe0001):

\(\bf hats=h\qquad scarves=s\\ \quad \\ 4s+6h=52\qquad then\qquad 2h=s+1\implies h=\cfrac{s+1}{2}\qquad so\\ \quad \\ 4s+6\left(\cfrac{s+1}{2}\right) = 52\)

OpenStudy (jdoe0001):

solve for "s" to find how much is a scarf

OpenStudy (anonymous):

Thank you so much @jdoe0001

OpenStudy (jdoe0001):

yw